# Neither Stochastic Parroting nor AGI: LLMs Solve Tasks through Context-Directed Extrapolation from Training Data Priors

### v1

@@ Abstract
In this position paper we raise critical awareness of a realistic view of LLM capabilities that eschews extreme alternative views that LLMs are either ‘stochastic parrots’ or in possession of ‘emergent’ advanced reasoning capabilities, which, due to their unpredictable emergence, constitute an existential threat. Our middle-ground view is that LLMs extrapolate from priors from their training data, and that a mechanism akin to in-context learning enables the targeting of the appropriate information from which to extrapolate. We call this ‘context-directed extrapolation.’ Under this view, substantiated though existing literature, while reasoning capabilities go well beyond stochastic parroting, such capabilities are predictable, controllable, not indicative of advanced reasoning akin to high-level cognitive capabilities in humans, and not infinitely scalable with additional training. As a result, fears of uncontrollable emergence of agency are allayed, while research advances are appropriately refocused on the processes of context-directed extrapolation and how this interacts with training data to produce valuable capabilities in LLMs. Future work can therefore explore alternative augmenting techniques that do not rely on inherent advanced reasoning in LLMs.
@@ 1 Introduction
Large Language Models (LLMs), primarily trained on the cloze-inspired task of next-token prediction, have demonstrated the ability to solve a range of tasks, some of which typically require reasoning in humans [10, 59, 49]. Understanding the basis for how LLMs achieve this is central to several key considerations, including the continued development of these models, their safety and security implications, and the contexts in which they should or should not be used [23]. However, despite, or perhaps because of, its importance, there has been little consensus on how LLMs are able to solve complex tasks. Currently, two predominant viewpoints exist: the first (alternative view 1) is that LLMs are essentially ‘stochastic parrots,’ predicting statistically likely outputs in line with their training objectives [6, 5, 34]. The second viewpoint (alternative view 2) posits that as LLMs are scaled up, in terms of both parameters and training data, they develop ‘emergent reasoning’ capabilities [10, 59, 49], which, in some cases, have been described as “sparks of AGI” (Artificial General Intelligence) [12].
@@
Neither of these viewpoints, we will argue, sufficiently explain the capabilities and limitations of state-of-the-art LLMs. Specifically, LLMs trained only on next token prediction can solve a range of non-memorizable tasks that require more than generating the next most likely token (see Section 3). Nonetheless, these models fail on a range of tasks which humans are able to trivially solve (Section 4), indicating that they are not capable of certain types of human-like reasoning. Note that our discussion of reasoning focuses on capabilities that are either already at the level of, or indicative of the potential for, advanced general-purpose reasoning akin to high level human cognitive capabilities. For brevity we will call this ‘advanced reasoning’ and we define levels of reasoning in Section 2.
@@
Despite the conflicting findings and viewpoints surrounding the explanation of LLM performance, there is one ability that LLMs possess indisputably: this is the ability of LLMs, trained solely on next-token prediction (referred to as Base LLMs), to solve a range of tasks when provided with a few examples in the input prompt—an ability known as in-context learning (ICL) or few-shot learning [11, 37]. For example, an LLM provided with the input prompt “1 + 3 = 5; 7 + 12 = 20; 8 + 3 =” will respond with ‘12.’ Notice that this requires the model to identify this task to be a form of modified addition ($a+b+1$) and perform the required modified calculation.We note that this alone serves as evidence against the stochastic parrot argument.
@@ 2 Defining Reasoning
In this research, we focus on the Cognitive Process dimension, which is an ordered set (from simple to more complex) of processes that subdivide reasoning: Remember, Understand, Apply, Analyze, Evaluate, Create. Within these processes, we focus on Remember, Understand and Apply. Our focus is guided by the fact that this is considered an ordered set of processes, such that the higher-level processes of Evaluate and Create rely upon mastery of the lower-level cognitive processes; within this classification we find evidence that LLMs are limited to Understand, which would preclude higher-order processes. [^1] We also caution the reader against assuming commonplace definitions of these terms—there are certainly levels of “understanding” that models do not achieve, and it is clear that models can do some types of evaluation and can create new texts, but this does not necessarily meet the definitional criteria of these terms as used in educational settings. Remember is defined as “Retrieving relevant knowledge from long-term memory,” and is split into subcategories of Recognizing and Recalling [29, p. 215]. Understand is defined as “Determining the meaning of instructional messages, including oral, written, and graphic communication,” and is split into finer grained categories of Interpreting, Exemplifying, Classifying, Summarizing, Inferring, Comparing, and Explaining (ibid.). Apply is defined as “Carrying out or using a procedure in a given situation,” but we also leverage Bloom’s original Application description, which involves solving a new, unseen problem by deploying and operationalizing previously learned information. Apply is split into Executing and Implementing.
@@ 3 Alternative View: Stochastic Parroting
The notion that LLMs are ‘Stochastic Parrots’ [^3] We note that stochastic parroting does not have a precise parallel in the literature on human cognitive processes described in Section 2. It relies strongly on Remembering, but rather than memory of chunks of information, it also involves extrapolation to guess the next most likely token, based upon what is in memory. Here, the term extrapolation refers to the mathematical, machine-learning sense: estimating values outside the range of known data points based on the observed relationship between variables. that respond with the next most likely token was first introduced by Bender et al. [6]. They rightly argue that statistical generalization from web data at scale can lead to the propagation of biases inherent to the statistical patterns present in the text on the web. We do not contradict this argument. However, Bender et al. [6] also suggests that LLMs merely generate the next most likely token. As demonstrated earlier using the modified arithmetic example, this claim is demonstrably false. The most likely token to follow the string ‘8 + 3’ is almost certainly ‘11’ and not, as in our example, ‘12.’ The fact that models output the correct answer when prompted with examples demonstrating an alternate task, which requires a modified interpretation of addition to mean $a+b+1$, provides clear evidence against the notion that these models are stochastic parrots.
@@
It is now undeniable that sufficiently large LLMs trained on sufficiently large corpora on the next-token prediction task exhibit in-context capabilities or the ability to solve tasks based on a few examples, as described above [11, 59, 50, 33, 62]. These capabilities, if nothing else, truly set LLMs apart and make LLMs fundamentally different from all previous machine learning systems.
@@ 4.2 Hallucinations
The second major piece of evidence against the notion of advanced reasoning in LLMs is their tendency to generate output that is not aligned with reality, a phenomenon known as hallucination [24]. This has also been referred to as confabulation and compared to human confabulation [52]. However, it is important to distinguish between the two, as there is no evidence to suggest that LLMs possess any agency (further discussed in Sections 5.1, 5.2.3 and 6), and because hallucinations can always be traced to the generation of a statistically most likely next token or pattern [22]. In fact, this second factor reinforces our perspective that LLMs are relying on priors from pre-training data guided by the context presented in the prompt. When the prompt lacks the necessary context to direct the model to the correct priors, LLM outputs default to the broader pre-training data (which determine the statistically most likely output), resulting in hallucinations.
@@ 5.1 Base Models: What Does Pre-Training Enable?
Olsson et al. [37] were the first to systematically explore the capabilities of LLMs in performing ICL with random tokens. They demonstrated that models can consistently complete patterns presented in their prompts, even when the patterns involve random tokens. Specifically, they showed that when presented with a sequence $[A][B]...[A]$, LLMs could consistently respond with $[B]$, even if $A$, $B$, and the intermediate tokens were all random. The use of random tokens in this experiment was intentional and significant. The fact that the sequences were entirely random suggests that the models were not relying on memorization. This provides compelling evidence against the notion that LLMs are merely stochastic parrots.
@@ 5.2.1 The Impact of Fine-Tuning
To clarify, we do not deny that LLMs’ ability to solve tasks increases as their scale grows, as argued by Kaplan et al. [26]. However, the ability of a model to understand, for example, social situations can be evaluated through tasks that involve questions about such scenarios. If the model performs above a random baseline on these tasks, we might describe this as an emergent ability. For instance, a model trained to answer questions about social situations at work that can then answer questions about social situations at home might exhibit generalization, but this alone would not qualify as ‘emergence’ of reasoning on the level that would imply advanced reasoning akin to human-like cognitive processes beyond the simplest levels of Remember and Understand. If the tasks are similar, the model’s performance is less indicative of emergent reasoning [7]. Understanding the degree to which instruction-tuned models can generalize to novel tasks is essential in assessing their true emergent abilities. Importantly, generalization over such a narrow domain is less indicative of the kind of high level cognitive capabilities that is of interest to us. Furthermore, recent research has demonstrated that, when the specific tasks used in instruction-tuning are taken into account, there is a statistically significant correlation between Base and instruction-tuned model performance [7] indicating that instruction tuning itself does not significantly alter the ability of LLMs.

### v2

@@ Abstract
In this position paper we raise critical awareness of a realistic view of LLM capabilities that eschews extreme alternative views that LLMs are either ‘stochastic parrots’ or in possession of ‘emergent’ advanced reasoning capabilities, which, due to their unpredictable emergence, constitute an existential threat. Our middle-ground view is that LLMs extrapolate from priors from their training data while using context to guide the model to the appropriate priors; we call this ‘context-directed extrapolation.’ Specifically, this context direction is achieved through examples in base models, leading to in-context learning, while instruction tuning allows LLMs to perform similarly based on prompts rather than explicit examples. Under this view, substantiated though existing literature, while reasoning capabilities go well beyond stochastic parroting, such capabilities are predictable, controllable, not indicative of advanced reasoning akin to high-level cognitive capabilities in humans, and not infinitely scalable with additional training. As a result, fears of uncontrollable emergence of agency are allayed, while research advances are appropriately refocused on the processes of context-directed extrapolation and how this interacts with training data to produce valuable capabilities in LLMs. Future work can therefore explore alternative augmenting techniques that do not rely on inherent advanced reasoning in LLMs.
@@ 1 Introduction
Large Language Models (LLMs), primarily trained on the cloze-inspired task of next-token prediction, have demonstrated the ability to solve a range of tasks, some of which typically require reasoning in humans [11, 62, 52]. Understanding the basis for how LLMs achieve this is central to several key considerations, including the continued development of these models, their safety and security implications, and the contexts in which they should or should not be used [25]. Simultaneously, better understanding the underlying framework used by LLMs to solve such tasks has significant implications for various research domains. For example, it would suggest domains wherein the use of LLMs would require additional guardrails and training.Therefore, providing a clear explanation for LLM performance is crucial. However, despite, or perhaps because of, its importance, there has been little consensus on how LLMs are able to solve complex tasks. Currently, two predominant viewpoints exist: the first (alternative view 1) is that LLMs are essentially ‘stochastic parrots,’ predicting statistically likely outputs in line with their training objectives [6, 5, 37]. The second viewpoint (alternative view 2) posits that as LLMs are scaled up, in terms of both parameters and training data, they develop ‘emergent reasoning’ capabilities [11, 62, 52], which, in some cases, have been described as “sparks of AGI” (Artificial General Intelligence) [13]. It is important to note that our discussion focuses on capabilities that are either already at the level of, or indicative of the potential for, advanced general-purpose reasoning akin to high level human cognitive capabilities. For brevity we will call this ‘advanced reasoning.’
@@
Neither of these viewpoints, we will argue, sufficiently explain the capabilities and limitations of state-of-the-art LLMs. Specifically, LLMs trained only on next token prediction can solve a range of non-memorizable tasks that require more than generating the next most likely token (see Section 3). Nonetheless, these models fail on a range of tasks which humans are able to trivially solve (Section 4), indicating that they are not capable of certain types of human-like reasoning. Note that our discussion of reasoning focuses on capabilities that are either already at the level of, or indicative of the potential for, advanced general-purpose reasoning akin to high level human cognitive capabilities.
@@
Despite the conflicting findings and viewpoints surrounding the explanation of LLM performance, there is one ability that LLMs possess indisputably: this is the ability of LLMs, trained solely on next-token prediction (referred to as Base LLMs), to solve a range of tasks when provided with a few examples in the input prompt—an ability known as in-context learning (ICL) or few-shot learning [12, 40]. For example, an LLM provided with the input prompt “1 + 3 = 5; 7 + 12 = 20; 8 + 3 =” will respond with ‘12.’ Notice that this requires the model to identify this task to be a form of modified addition ($a+b+1$) and perform the required modified calculation. [^1] We test this on LLaMA 3 70B Base
The idea that LLMs are just “stochastic parrots” is challenged by their ability to learn in-context. For example, research by Bonial et al. [10] shows that LLMs can be taught entirely novel phrases, or Multi-Word Expressions (MWEs), on the fly. After seeing just one example defining “winking at pringles,” a novel MWE assigned the meaning ‘to indulge in frivolity,’ models could correctly interpret and even reason about new uses of the phrase. Since these novel MWEs could not have been in the training data, this performance demonstrates a capability beyond simply predicting the next most likely word.
@@ 2 Defining Reasoning
In this research, we focus on the Cognitive Process dimension, which is an ordered set (from simple to more complex) of processes that subdivide reasoning: Remember, Understand, Apply, Analyze, Evaluate, Create. Within these processes, we focus on Remember, Understand and Apply. Our focus is guided by the fact that this is considered an ordered set of processes, such that the higher-level processes of Evaluate and Create rely upon mastery of the lower-level cognitive processes; within this classification we find evidence that LLMs are limited to Understand, which would preclude higher-order processes. [^2] We also caution the reader against assuming commonplace definitions of these terms—there are certainly levels of “understanding” that models do not achieve, and it is clear that models can do some types of evaluation and can create new texts, but this does not necessarily meet the definitional criteria of these terms as used in educational settings. Remember is defined as “Retrieving relevant knowledge from long-term memory,” and is split into subcategories of Recognizing and Recalling [31, p. 215]. Understand is defined as “Determining the meaning of instructional messages, including oral, written, and graphic communication,” and is split into finer grained categories of Interpreting, Exemplifying, Classifying, Summarizing, Inferring, Comparing, and Explaining (ibid.). Apply is defined as “Carrying out or using a procedure in a given situation,” but we also leverage Bloom’s original Application description, which involves solving a new, unseen problem by deploying and operationalizing previously learned information. Apply is split into Executing and Implementing.
@@ 3 Alternative View: Stochastic Parroting
The notion that LLMs are ‘Stochastic Parrots’ [^4] We note that stochastic parroting does not have a precise parallel in the literature on human cognitive processes described in Section 2. It relies strongly on Remembering, but rather than memory of chunks of information, it also involves extrapolation to guess the next most likely token, based upon what is in memory. Here, the term extrapolation refers to the mathematical, machine-learning sense: estimating values outside the range of known data points based on the observed relationship between variables. that respond with the next most likely token was first introduced by Bender et al. [6]. They rightly argue that statistical generalization from web data at scale can lead to the propagation of biases inherent to the statistical patterns present in the text on the web. We do not contradict this argument. However, Bender et al. [6] also suggests that LLMs merely generate the next most likely token.
As shown by the novel MWE example in Section 1, this claim is demonstrably false. In the novel MWE interpretation task of Bonial et al. [10], a stochastic parrot generating statistically probable outputs would have no existing knowledge of the assigned phrasal meaning of “winking at pringles.” Rather, it would only have access to information about the meanings of the individual lexical items as used in pretraining data. It would therefore likely produce a literal or nonsensical completion. The fact that models can use the definition provided in the prompt to ‘learn’ the MWE’s novel meaning (to indulge in frivolity) and correctly reason about it in new situations provides clear evidence against the notion that they are just stochastic parrots.
@@
Furthermore, it is now undeniable that sufficiently large LLMs trained on sufficiently large corpora on the next-token prediction task exhibit in-context capabilities or the ability to solve tasks based on a few examples, as described above [12, 62, 53, 36, 65].
For example, while LLMs trained solely on next-token prediction (i.e. base models) cannot solve tasks that require abstract reasoning, they can solve such problems when provided with examples through ICL [36]. An illustration of this is the following logical deduction problem from the Big-Bench benchmark:
Question: On a shelf, there are five books: a gray book, a red book, a purple book, a blue book, and a black book. The red book is to the right of the gray book. The black book is to the left of the blue book. The blue book is to the left of the gray book. The purple book is the second from the right.
Targets: ‘The gray book is the leftmost.’: 0; ‘The red book is the leftmost.’: 0; ‘The purple book is the leftmost.’: 0; ‘The blue book is the leftmost.’: 0; ‘The black book is the leftmost.’: 1
The difference in performance when base models are provided with ICL examples and when they are not offers clear evidence for our framework of context-directed extrapolation. Base models fail on such reasoning tasks because they require explicit examples in the prompt to provide the ‘context’ needed to guide their extrapolation. In contrast, instruction-tuned models can derive this context directly from the prompt, allowing them to solve the problem without needing any examples, as suggested by Lu et al. [36].
Finally, the argument against ‘stochastic parroting’ becomes even stronger when you use novel words, which forces a model to reason instead of just relying on statistical patterns. In addition to the novel MWE experiments presented above, the Winodict benchmark [20] provides a similar test case by adapting the Winograd Schema Challenge [32], which offers a classic method for evaluating a system’s capacity for common-sense reasoning through the task of anaphora resolution. The original Winograd task consists of sentence pairs that differ by only a single word, a change that completely alters the referent of an ambiguous pronoun. For example, consider ‘The man could not lift his son because he was so weak.’ and ‘The man could not lift his son because he was so heavy.’ In the first sentence, ‘he’ refers to the man and in the second, ‘he’ refers to the son.
The Winodict benchmark replaces the critical content word with a "nonce" word whose definition is provided entirely within the prompt. This transforms the task from a test of stored world knowledge to one of dynamic, in-context semantic application. For instance, a model is presented with the following: The verb ‘to plest’ means to be scared of. The city councilmen refused the demonstrators a permit because they plested violence. Here, the resolution of the pronoun ‘they’ is entirely dependent on the model’s ability to parse the ad-hoc definition of ‘plest’ and apply it to the sentence’s causal structure.
These dynamic, in-context capabilities truly set LLMs apart and make LLMs fundamentally different from all previous machine learning systems and challenge the notion of them being stochastic parrots.
@@ 4.2 Hallucinations
The second major piece of evidence against the notion of advanced reasoning in LLMs is their tendency to generate output that is not aligned with reality, a phenomenon known as hallucination [26]. This has also been referred to as confabulation and compared to human confabulation [55]. However, it is important to distinguish between the two, as there is no evidence to suggest that LLMs possess any agency (further discussed in Sections 5.1, 5.2.3 and 7), and because hallucinations can always be traced to the generation of a statistically most likely next token or pattern [24]. In fact, this second factor reinforces our perspective that LLMs are relying on priors from pre-training data guided by the context presented in the prompt. When the prompt lacks the necessary context to direct the model to the correct priors, LLM outputs default to the broader pre-training data (which determine the statistically most likely output), resulting in hallucinations.
@@ 5.1 Base Models: What Does Pre-Training Enable?
Olsson et al. [40] were the first to systematically explore the capabilities of LLMs in performing ICL with random tokens. They demonstrated that models can consistently complete patterns presented in their prompts, even when the patterns involve random tokens. Specifically, they showed that when presented with a sequence [A][B][A][A][B]...[A], LLMs could consistently respond with $[B]$, even if $A$, $B$, and the intermediate tokens were all random. The use of random tokens in this experiment was intentional and significant. The fact that the sequences were entirely random suggests that the models were not relying on memorization. This provides compelling evidence against the notion that LLMs are merely stochastic parrots.
@@ 5.2.1 The Impact of Fine-Tuning
To clarify, we do not deny that LLMs’ ability to solve tasks increases as their scale grows, as argued by Kaplan et al. [28]. However, the ability of a model to understand, for example, social situations can be evaluated through tasks that involve questions about such scenarios. If the model performs above a random baseline on these tasks, we might describe this as an emergent ability. For instance, a model trained to answer questions about social situations at work that can then answer questions about social situations at home might exhibit generalization, but this alone would not qualify as ‘emergence’ of reasoning on the level that would imply advanced reasoning akin to human-like cognitive processes beyond the simplest levels of Remember and Understand. However, if a model is explicitly trained to answer questions about social situations, it may be able to apply this knowledge within a narrow domain, but generalizing to novel situations without clear out-of-domain performance would not constitute emergent reasoning. If the tasks are similar, the model’s performance is less indicative of emergent reasoning [7]. Understanding the degree to which instruction-tuned models can generalize to novel tasks is essential in assessing their true emergent abilities. Importantly, generalization over such a narrow domain is less indicative of the kind of high level cognitive capabilities that is of interest to us. Furthermore, recent research has demonstrated that, when the specific tasks used in instruction-tuning are taken into account, there is a statistically significant correlation between Base and instruction-tuned model performance [7] indicating that instruction tuning itself does not significantly alter the ability of LLMs.
@@ 5.2.3 Explaining Emergent Abilities through Contextual-Extrapolation
6 Context Directed Extrapolation vs Stochastic Parroting
We would like to acknowledge that Context Directed Extrapolation and Stochastic Parroting share some commonalities. Therefore, this section clarifies the commonalities and differences between the stochastic parroting perspective and the perspective of context-directed extrapolation from training data priors.
Commonalities. From the perspective of base model performance, there is no functional difference between context-directed extrapolation and stochastic parroting. As previously discussed, base models consistently fail on tasks requiring reasoning when they are presented without examples. One can argue that the examples simply form a long context, making the correct answer the most probable sequence completion. This makes both theories appear to describe the same mechanism: the model completes a given sequence based on statistical patterns. Consequently, when analyzing base models alone, the two views are indistinguishable.
Differences. The functional divergence between these two views becomes apparent with instruction-tuned models. While a base model’s output can be seen as the most probable completion of a long sequence of examples, the context-directed extrapolation framework posits that instruction-tuning enables a different mechanism. Our framework suggests that instruction-tuning allows a model to interpret an instruction as a directive to construct an implicit task context, rather than as a literal sequence to be continued. This context activates relevant priors from the model’s pre-training data, enabling it to solve the specified task. The critical evidence supporting this distinction is that instruction-tuned models can solve logical deduction problems without any examples. This phenomenon is directly accounted for by context-directed extrapolation but cannot be explained by a strict stochastic parroting model. This functional gap becomes even more stark in tasks involving novel words, as described previously.
Grounding An important implication of context-directed extrapolation is that it facilitates a limited form of grounding. This is not grounding in the sense of connecting language to embodied experience, but rather a mechanism by which the model can access information that is not explicitly present in the surface text. Because the process involves extrapolating from priors activated by the prompt, a model can, for instance, take the definition of a nonce verb and productively apply that new meaning in novel contexts.
Therefore, while there is some overlap, our context-directed perspective is fundamentally distinct from a stochastic parroting mechanism, which by definition is incapable of handling novel inputs that fall outside its memorized distribution. The demonstrated ability of LLMs to extend prompt-based definitions, apply abstract patterns, and generate context-appropriate interpretations indicates that extrapolation provides access to information not reducible to surface statistics. In this sense, context-directed extrapolation offers a framework to explaining LLM capabilities and limitations, including a form of limited grounding, which remain contingent upon the model’s training priors and the context supplied at inference.